40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15047     Entry time: Mon Nov 25 22:10:26 2019
Author: shruti 
Type: Update 
Category: NoiseBudget 
Subject: Diagnostics  

This is to help troubleshoot the excess noise measured earlier.

The following channels were measured at GPS times 1258586880 s and 1258597457 s, corresponding to low and high Power Recycling Gain (PRG) respectively.

Excess noise was seen between 25-110 Hz in the high PRG case when compared to the low PRG case in the following channels:

 

C1:LSC-CARM-IN1_DQ (shown in Attachment 1 where the reference is low PRG)

C1:ALS-Y_ERR_MON_OUT_DQ

C1:ALS-BEAT{X,Y}_FINE_PHASE_OUT_DQ

C1:SUS-ETM{X,Y} _SENSOR_{LL,LR,UL,UR}

C1:ALS-TRX_OUT_DQ

 

Surprisingly, it was also seen to a smaller extent in (refer Attachment 3)

C1:SUS-ITMX_SENSOR_{LL,LR,UL,UR} 

 

A different type of noise spectrum, attributed to known electronic effects, was observed for

C1:SUS-ITMY_SENSOR_{LL,UL}    (refer Attachment 2)

 

These did not show any significant change in the noise spectrum:

C1:LSC-DARM-IN1_DQ (shown in Attachment 1 where the reference is low PRG)

C1:ALS-X_ERR_MON_OUT_DQ

C1:ALS-TRY_OUT_DQ

C1:SUS-ITMY_SENSOR_{LL,LR,UL,UR} 

C1:SUS-ITMY_SENSOR_{LR,UR}  (refer Attachment 2)

 

Broadband noise in:

C1:LSC-PO{X,Y}11_I_ERR_DQ

 

Attachment 1: LSC.pdf  42 kB  | Hide | Hide all
LSC.pdf
Attachment 2: ITMY.pdf  74 kB  | Hide | Hide all
ITMY.pdf
Attachment 3: ITMX_L.pdf  41 kB  | Hide | Hide all
ITMX_L.pdf
ELOG V3.1.3-